OpenStudy (anonymous):

At a certain time of day, a person 6 ft tall casts a 4 ft shadow. How long is the shadow cast by a 21 ft tree at the same time? A. 14 ft B. 15 ft C. 12 ft D. 19 ft

OpenStudy (zzr0ck3r):

so we have a ratio of 6::4 we want to keep that ratio so 21::? = 6::4 \[\frac{21}{x}=\frac{6}{4}\\so\\x=\frac{4*21}{6}=?\]

OpenStudy (anonymous):

@ ; zzr0ck3r : I think you re fall ... I have a better way >

OpenStudy (anonymous):

21/x - 6/4 ? Is it true ?

OpenStudy (anonymous):

ft tall 6 21 ft show 4 x 4.21/6=84/6=42/3=14

OpenStudy (zzr0ck3r):

I think you mean 4*21/6 not 4.21/6 good job, this is the exact same thing I did, but looks very ugly. you right....@E.ali

OpenStudy (goformit100):

"Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."